At 30,000 feet above sea level, atmospheric pressure is only 224 torr. That’s why airplane cabins are pressurized. There isn’t enough air for you to breathe outside the plane! If we took a balloon filled to 3.86L at room temperature (21°C) and 0.947 atm, what would its volume be at 30,000 feet above sea level (outside the airplane!) if the temperature also reduced to -35.°C?
Answer #1

ANSWER

According to question Given that,

See level atmospheric pressure (P1) = 224 torr, And temperature (T1) = -35°C = -35 + 273 = 238K And balloon filled to V2 = 3.86L and P2 = 0.947 atm= 760 × 0.947= 719.72 torr. And T2 = 21°C = 21+273 = 294K

According to Combined ideas gas equation,

P1V1/T1 = P2V2/T2

V1 = P2V2T1/T2P1 = 719.72 × 3.86 × 238/294×224 = 10.03

V1 = 10.03L

Volume will be 10 L at 30,000 see level if the temperature also reduced to -35°C.